OpenStudy (anonymous):

what is the ratio of 3:11

ganeshie8 (ganeshie8):

\(\large 3 : 11\) is same as \(\large \frac{3}{11}\)

OpenStudy (anonymous):

15:55 because 3*5 = 15 and 11*5 = 55.